QUESTION 2 T(n) = Tỉn/2) + n T(n) = Theta QUESTION 3 T(n) = 2 Tỉn/2) + 1 T(n) = Theta QUESTION 4 T(n) = 4 Tỉn/2) + n^2 T(
Answer #1

Page No. Date: / Solution 02 Well solve this using Masters method T(n) = at(n/6) + f(n), az1, 671 Ten) = T(1/2) + W a=1, 6=

Please give an upvote if you liked my solution. Thank you:)